# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

### v4

@@ Abstract
Feedback optimization algorithms compute inputs to a system using real-time output measurements, which helps mitigate the effects of disturbances. However, existing work often models both system dynamics and computations in either discrete or continuous time, which may not accurately model some applications. In this work, we model linear system dynamics in continuous time, and we model the computations of inputs in discrete time. Therefore, we present a novel hybrid systems model of feedback optimization. We first establish the well-posedness of this hybrid model and establish completeness of solutions while ruling out Zeno behavior. Then we show the state of the system converges exponentially fast to a ball of known radius about a desired goal state. Next we analytically show that this system is robust to perturbations in (i) the values of measured outputs, (ii) the matrices that model the linear time-invariant system, and (iii) the times at which inputs are applied to the system. Simulation results confirm that this approach successfully mitigates the effects of disturbances.
@@ I Introduction
We seek to show that feedback optimization retains its robustness guarantees with continuous-time dynamics and discrete-time optimization, and we therefore develop a hybrid systems model of it. Related work in [15, 16, 17] considers a continuous-time system and discrete-time computations in a sampled-data setting, though we develop new analytical robustness guarantees by developing a hybrid model in the framework of [18]. We discuss this point more below.
@@
  • We analytically prove the robustness of the hybrid model to perturbations by showing that there is bounded difference between the solutions to a perturbed system and the solutions to a nominal system (Theorem 3).
  • We show in simulation that hybrid feedback optimization successfully rejects disturbances (Section VI).
Prior work in [15, 16, 17] combines continuous-time dynamics and discrete-time optimization in a sampled-data feedback optimization setting. The main results in [15, 16] show that the closed-loop system is stable for large enough sample times and that it is practically stable under time-varying disturbances with bounded variation. Results in [17] show global exponential stability of the closed-loop system when inputs to the system change at a fixed rate.
We differ from those prior works by considering a hybrid systems model in the framework of [18], which unlocks robustness analyses that we apply in this work. We use this model to show that hybrid feedback optimization is simultaneously robust to several types of disturbances, including disturbances in the values of measured outputs, errors in the times at which new inputs are applied to the system, and errors in all of the matrices that model the linear time-invariant system that is being controlled. To the best of our knowledge, this paper is the first to analytically prove that a hybrid/sampled-data feedback optimization model is simultaneously robust to all of these disturbances.
The rest of the paper is organized as follows. Section II provides background. Section III gives problem statements and the hybrid feedback optimization model. Section IV derives properties of its solutions. Then Section V establishes the solutions’ convergence and the system’s robustness. Section VI presents simulations, and Section VII concludes.
@@ II-A Notation
Let $\mathbb{R}$ denote the reals and let $\mathbb{N}$ denote the non-negative integers. For a differentiable function $\Phi:\mathbb{R}^{m}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$, let $\nabla_{u{}}\Phi$ denote the partial derivative with respect to its first argument. The symbol $I_{b}$ denotes the identity matrix of dimension $b$. The $2$-norm of a vector $x$ is denoted $\left\|x\right\|$. For a non-empty, compact, convex set $\mathcal{Z}$, the symbol $\Pi_{\mathcal{Z}}[v]$ denotes the Euclidean projection of a point $v$ onto $\mathcal{Z}$, i.e., $\Pi_{\mathcal{Z}}[v]=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\|v-z\|$. We use $\lambda_{i}(N)$ to denote the $i^{th}$ eigenvalue of a matrix $N$, and we use $\mathfrak{Re}\{a\}$ to denote the real part of a complex number $a$. We also write $\lambda_{min}(N)$ and $\lambda_{max}(N)$ for the smallest and largest (real) eigenvalues of a symmetric matrix $N$, respectively. Given $r\geq 0$, we use $B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}):=\{x\in\mathbb{R}^{n}:\|x-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\|\leq r\}$ to denote the closed Euclidean ball of radius $r$ about the point $\tilde{x}\in\mathbb{R}^{n}$. For a set $S$ and a point $x$, we use $\|x\|_{S}:=\inf_{s\in S}\|x-s\|$.
@@ II-B Feedback Optimization Background and Setup
Suppose we have the linear time invariant (LTI) system
@@
where $x\in\mathbb{R}^{n}$ is the system’s state, $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is its input, $\mathcal{U}$ is a non-empty, compact, convex set, and $y\in\mathbb{R}^{p}$ is its output. The vector $d\in\mathbb{R}^{p}$ is a constant, unknown disturbance (e.g., bias), which is a typical component of feedback optimization problem formulations [1, 10, 24, 9, 13]. Such disturbances arise for example in various power systems applications [1, 24, 9, 10, 13].
@@
We adopt the “stabilize then optimize” approach [15] in which we suppose that a stabilizing controller has already been applied to the system. If Assumption 1 is not satisfied, then it can be enforced for any stabilizable system by doing pre-feedback with a stabilizing controller.
Assumption 1 ensures that a system will eventually reach steady state, which allows its steady-state behavior to be optimized. From (1) the steady-state input-to-output map is $u\mapsto Hu+d$, where $H:=-\Psi A^{-1}B$.
@@
While one could envision using an estimator to determine $d$, we are interested in cases in which the matrices $A$, $B$, and $\Psi$ are not known exactly, and an observer may have poor accuracy under these conditions. We also emphasize that knowledge of $H:=-\Psi A^{-1}B$ is used only for analysis and is not required to implement our algorithm.
@@
$\min_{u,y{}}\quad$ $\Phi(u,y{})$ (2a)
subject to $y{}=Hu+d,\,\,u\in\mathcal{U},\,\,y{}\in\mathbb{R}^{p},$ (2b)
where $\mathcal{U}$ is a non-empty, compact, convex set and $\Phi:\mathbb{R}^{m}\times\mathbb{R}^{p}\to\mathbb{R}$ is strongly convex in $(u,y)$. These properties ensure (2a)-(2b) has a unique solution.
@@
$\min_{u}\quad$ $\tilde{\Phi}(u):=\Phi(u,Hu+d)$ (3)
subject to $u\in\mathcal{U}.$ (4)
However, the problem in (3) cannot be solved in practice because the substitution $y=Hu+d$ would require exact knowledge of the disturbance $d$, which may not be available. Instead, feedback optimization is used to repeatedly measure $y$ (which is perturbed by $d$) and then optimize over $u$. In this work, $y$ is sampled at discrete instants of time, and we use $y_{s}$ to denote its sampled value.
@@
The underlying LTI system need not always be at steady state, but we will use a standard technique in the feedback optimization literature to approximate its outputs as coming from a system at steady state [8, 9, 10]. Mathematically, we approximate a sampled output $y_{s}$ as coming from the steady-state map $y_{s}=Hu+d$ when the input to the system is $u$. This approximation is justified, for example, when the dynamics of the system converge sufficiently quickly. Our results in Sections IV and V still analyze the dynamics for the system, and they use this approximation only to relate outputs to inputs.
We study a closed-loop system that connects the LTI system with a gradient descent algorithm that computes the input $u$. With a sampled output $y_{s}$, the optimization update law is $u_{k+1}=\Pi_{\mathcal{U}}\big[u_{k}-\gamma\nabla_{u}\Phi(u_{k},y_{s})\big]$, where $u_{k}$ is the $k^{th}$ iterate of the gradient descent algorithm.
Then the closed-loop interconnected LTI system is
@@ II-C Background on Hybrid Systems
where $\zeta\in\mathbb{R}^{n}$ is the system’s state vector and the maps $F$ and $G$ are set valued in general. The function $F$ defines the flow map and governs the continuous dynamics within the flow set $C$, while $G$ defines the jump map, which models the system’s discrete behavior within the jump set $D$.
@@
If a hybrid system satisfies the hybrid basic conditions, then it is well-posed by [18, Theorem 6.30]. We use this property in Section V to show that errors in the models of $F$ and $G$ up to a certain threshold produce bounded changes in the resulting closed-loop system trajectories over compact time horizons. It is not automatic to formulate a well-posed hybrid system, and this paper will do so for a hybrid model of feedback optimization.
@@ III A Hybrid Model for Feedback Optimization
Prove the convergence of solutions toward a desired goal state and bound their steady-state error.
@@ III-B Optimization Problem Setup
$\Phi(u,y{})=\frac{1}{2}u^{\top}Q_{u{}}u+\frac{1}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(y_{s}-\hat{y})}^{\top}Q_{y{}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(y_{s}-\hat{y})}$, (8)
@@ III-D Flow Map Definition
$\dot{\zeta}=F(\zeta):=\left(\begin{array}[]{c}Ax+Bu\\ 0\\ 0\\ 0\\ -\hskip-2.0pt1\\ -\hskip-2.0pt1\end{array}\right).$ (12)
@@ III-E Jump Map Definition
$G_{1}(\zeta)=\left(\begin{array}[]{c}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\big[z-\gamma\nabla_{u}\Phi(z,y_{s})\big]\\ \tau_{c}\\ \tau_{g,comp}\end{array}\right).$ (13)
In case (ii), when $\tau_{c}=0$ and $\tau_{g}>0$, a new input is applied to the system, a new output is sampled, and the timer $\tau_{c}$ resets to some point in the interval $\tau_{c,min},\tau_{c,max}]$, where $0<\tau_{c,min}\leq\tau_{c,max}$. This range of times represents indeterminacy in the amount of time that elapses between the application of successive inputs to the system.
Assumption 2.
Assumption 2 ensures that, when the system is properly initialized, at least $\ell$ gradient descent iterations are performed between any consecutive changes in the input. This condition is a mild form of timescale separation and it will be used later in Theorem 1. When the input changes, it is set equal to $z$. When $\tau_{c}$ reaches 0, the LTI system output is sampled and stored in $y_{s}$, which is held constant until the next sample. The states $x$ and $z$ do not change, and the jump map is
$G_{2}(\zeta)=\left(\begin{array}[]{c}x\\ z\\ Hu+d\\ z\\ {[\tau_{c,min},\tau_{c,max}]}\\ \tau_{g}\end{array}\right),$ (14)
where, as described in Remark 3, we approximate the output $y_{s}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Psi}x+d$ with $y_{s}=Hu+d$.
In case (iii), where both $\tau_{c}=0$ and $\tau_{g}=0$, we combine cases (i) and (ii), and the system executes $G_{1}$ and then $G_{2}$ or $G_{2}$ and then $G_{1}$. The full jump map $G$ is defined as
@@
where $G_{3}(\zeta)=G_{1}(\zeta)\cup G_{2}(\zeta)$. The jump set $D$ is equal to
$D=D_{1}\cup D_{2},$ (16)
where $D_{1}=\{\zeta\in\mathcal{X}:\tau_{g}=0\}$ and $D_{2}=\{\zeta\in\mathcal{X}:\tau_{c}=0\}$.
The full hybrid model of feedback optimization is
@@ IV-B Algorithm Framework
Consider an initial condition $\nu$ that satisfies
$\tau_{c}(0,0)\in[\tau_{c\min},\tau_{c,\max}]\textnormal{ and }\tau_{g}(0,0)=\tau_{g,comp},$ (20)
and consider a solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$. We use $\alpha(i)$ to denote the number of gradient descent iterations that are performed when computing the $(i+1)^{th}$ value of the input $u$. Assumption 2 implies that $\alpha(i)\geq\ell\geq 1$ for all $i\in\mathbb{N}$.
After the initial input to the system $u(0,0)$ is applied, the system performs $\alpha(0)$ gradient descent iterations (which are $\alpha(0)$ Case (i) jumps) before the input to the LTI system is changed. The initial condition for these iterates is $z_{0}(0,0)=u(0,0)$, and the $k^{th}$ such iterate is denoted $_{k}(t_{k},k); zz$ without a subscript always denotes a state of $\mathcal{H}_{FO}$, and subscripts are used only in our analyses to keep track of how many iterations have been completed when computing the next input to the LTI system. Before $u$ jumps for the first time, $\alpha(0)$ iterates are computed, the last of which is denoted $z_{\alpha(0)}\big(t_{\alpha(0)},\alpha(0)\big)$.
@@
$u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)=\\ z_{\alpha(p-1)}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha(p-1)+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha(p-1)+p-1),$ (23)
which indicates that the new value of the input is equal to the most recently computed optimization iterate. The initial iterate for computing the next input is $z_{0}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)$, and it is set equal to the same value as $u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)$.

### v5

@@ Abstract
Feedback optimization algorithms compute inputs to a system using real-time output measurements, which helps mitigate the effects of disturbances. However, existing work often models both system dynamics and computations in either discrete or continuous time, which may not accurately model some applications. In this work, we model linear system dynamics in continuous time and the computation of inputs in discrete time, leading to a new hybrid system model of feedback optimization. First, we establish well-posedness of this hybrid model and establish completeness of solutions while ruling out Zeno behavior. Then, we show that the state of the system converges exponentially fast to a ball of known radius about a desired goal state. Next, we analytically show that this system is robust to perturbations over bounded (hybrid) time horizons in (i) the values of measured outputs, (ii) the matrices that model the linear time-invariant system, and (iii) the times at which inputs are applied to the system. Simulation results confirm that this approach successfully mitigates the effects of disturbances.
@@ I Introduction
We seek to show that feedback optimization retains its robustness guarantees with continuous-time dynamics and discrete-time optimization, and we therefore develop a hybrid systems model of it.
@@
  • We prove robustness of the hybrid model to perturbations by showing that over bounded (hybrid) time horizons there is bounded difference between the solutions to a perturbed system and the solutions to a nominal system (Theorem 3).
  • We validate with simulations that hybrid feedback optimization successfully rejects disturbances (Section VI).
Prior work in [3, 7, 5] combines continuous-time dynamics and discrete-time optimization in a sampled-data feedback optimization setting, though we develop new analytical robustness guarantees by developing a hybrid model in the framework of [10]. The main results in [3, 7] show that the closed-loop system is stable for large enough sample times and that it is practically stable under time-varying disturbances with bounded variation. Results in [5] show global exponential stability of the closed-loop system when inputs to the system change at a fixed rate (with periodic sampling times).
Our results consider a hybrid system model in the framework of [10], which allows us to characterize system behavior at all times (permits aperiodic sampling), and unlocks robustness analyses that we apply in this work. We use this model to show that hybrid feedback optimization is simultaneously robust to several types of disturbances, including disturbances in the values of measured outputs, errors in the times at which new inputs are applied to the system, and errors in all of the matrices that model the linear time-invariant system that is being controlled. To the best of our knowledge, this paper is the first to analytically prove that a hybrid/sampled-data feedback optimization model is simultaneously robust to all of these disturbances.
The rest of the paper is organized as follows. Section II provides background. Section III gives problem statements and the hybrid feedback optimization model. Section IV derives properties of its solutions. Then, Section V establishes the solutions’ convergence and the system’s robustness. Section VI presents simulations. Due to space constraints, proofs will be published elsewhere.
@@ II-A Notation
Let $\mathbb{R}$ denote the reals and let $\mathbb{N}$ denote the non-negative integers. For a differentiable function $\Phi\!:\!\mathbb{R}^{m}\!\times\!\mathbb{R}^{p}\!\rightarrow\!\mathbb{R}$, let $\nabla_{u{}}\Phi$ denote the partial derivative with respect to its first argument. The symbol $I_{b}$ denotes the identity matrix of dimension $b$. The $2$-norm of a vector $x$ is denoted $\left\|x\right\|$. For a non-empty, compact, convex set $\mathcal{Z}$, the symbol $\Pi_{\mathcal{Z}}[v]$ denotes the Euclidean projection of a point $v$ onto $\mathcal{Z}$, i.e., $\Pi_{\mathcal{Z}}[v]=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\|v-z\|$. We denote the diameter of a non-empty, compact, convex set $\mathcal{Z}$ by $d_{\mathcal{Z}}$. We use $\lambda_{i}(N)$ to denote the $i^{th}$ eigenvalue of a matrix $N$, and we use $\mathfrak{Re}\{a\}$ to denote the real part of a complex number $a$. We also write $\lambda_{min}(N)$ and $\lambda_{max}(N)$ for the smallest and largest (real) eigenvalues of a symmetric matrix $N$, respectively. Given $r\!\geq\!0$, we use $B_{r}(\tilde{x})\!:=\!\{x\!\in\!\mathbb{R}^{n}\!:\!\|x-\tilde{x}\|\!\leq\!r\}$ to denote the closed Euclidean ball of radius $r$ about the point $\tilde{x}\in\mathbb{R}^{n}$. For a set $S$ and a point $x$, we use $\|x\|_{S}\!:=\!\inf_{s\in S}\!\|x-s\|$.
@@ II-B Feedback Optimization Background and Setup
Suppose we have the linear time-invariant (LTI) system
@@
where $x\in\mathbb{R}^{n}$ is the system’s state, $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is its input, $\mathcal{U}$ is a non-empty, compact, convex set, and $y\in\mathbb{R}^{p}$ is its output. The vector $d\in\mathbb{R}^{p}$ is a constant, unknown disturbance (e.g., bias), which is a typical component of feedback optimization problem formulations [12, 27, 6, 11, 28]. Such disturbances arise for example, in various power systems applications [12, 6, 11, 27, 28].
@@
Without loss of generality, the eigenvalues of $A$ are ordered such that $\lambda_{n}\leq\lambda_{n-1}\leq\cdots\leq\lambda_{1}<0$.
@@
We adopt the “stabilize then optimize” approach [3, 27] in which we suppose that a stabilizing controller has already been applied to the system. If Assumption 1 is not satisfied, then it can be enforced for any stabilizable system by doing pre-feedback with a stabilizing controller.
Assumption 1 ensures that (1) will eventually reach steady state when its input is constant, which allows its steady-state behavior to be optimized. From (1), the steady-state input-to-output map is $u\mapsto Hu+d$, where $H:=-\Psi A^{-1}B$.
@@
While one could envision using an estimator to determine $d$, we are interested in cases in which the matrices $A$, $B$, and $\Psi$ are not known exactly, and an observer may have poor accuracy under these conditions. For implementation, exact knowledge of $H$ is unnecessary as errors are allowed as covered in Section V-D.
@@
[…]
where $\Phi:\mathbb{R}^{m}\times\mathbb{R}^{p}\to\mathbb{R}$ is strongly convex in $(u,y)$. These properties ensure (2a)-(2b) has a unique solution.
@@
$\underset{u}{\operatorname{minimize}}\quad\tilde{\Phi}(u,Hu+d)\quad\operatorname{subject\penalty 10000\ to}\quad u\in\mathcal{U}$ (3)
However, the problem in (3) cannot be solved in practice because the substitution $y=Hu+d$ would require exact knowledge of the disturbance $d$, which may not be available. Instead, feedback optimization is used to repeatedly measure $y$ and then optimize over $u$. In this work, $y$ is sampled at discrete instants of time, and we use $y_{s}$ to denote its sampled value.
@@
The underlying LTI system need not always be at steady state, but we will use a standard technique in the feedback optimization literature to approximate its outputs as coming from a system at steady state [20, 11, 27]. Mathematically, we approximate a sampled output $y_{s}$ as coming from the steady-state map $y_{s}=Hu+d$ when the input to the system is $u$. This approximation is justified, for example, when the dynamics of the system converge sufficiently quickly. Our results in Sections IV and V still analyze the dynamics for the system, and they use this approximation only to relate outputs to inputs. Since we optimize over $u$, we must account for $u$ when deriving the gradient; however, in its computation we use the sampled value instead.
We study a closed-loop system that connects the LTI system with a gradient descent algorithm that computes the input $u$. With a sampled output $y_{s}$, the optimization update law is $u_{k+1}=\Pi_{\mathcal{U}}\big[u_{k}-\gamma\nabla_{u}\Phi(u_{k},y_{s})\big]$, where $u_{k}$ is the $k^{th}$ iterate of the gradient descent algorithm. Then, the closed-loop interconnected LTI system is
@@ II-C Background on Hybrid Systems
where $\zeta\in\mathbb{R}^{n}$ is the system’s state vector and the maps $F$ and $G$ are set valued in general. The set-valued map $F$ defines the flow map and governs the continuous dynamics within the flow set $C$, while $G$ defines the jump map, which models the system’s discrete behavior within the jump set $D$.
@@
If a hybrid system satisfies the hybrid basic conditions, then it is well-posed by [10, Theorem 6.30]. We use this property in Section V to show that errors in the models of $F$ and $G$ up to a certain threshold produce bounded changes in the resulting closed-loop system trajectories over compact time horizons. It is not automatic to formulate a well-posed hybrid system, and this paper will do so for a hybrid model of feedback optimization.
@@ III A Hybrid Model for Feedback Optimization
Bound the steady-state error of the system relative to a desired goal state in terms of system parameters.
@@ III-B Optimization Problem Setup
[…]
@@ III-D Flow Map Definition
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F(\zeta)}:=\left(\begin{array}[]{c}Ax+Bu\\ 0\\ 0\\ 0\\ -\hskip-2.0pt1\\ -\hskip-2.0pt1\end{array}\right)\quad\textnormal{ for all }\zeta\in C.$ (11)
@@ III-E Jump Map Definition
$G_{1}(\zeta)=\left(\begin{array}[]{c}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\big[z-\gamma\nabla_{u}\Phi(z,y_{s})\big]\\ \tau_{c}\\ \tau_{g,comp}\end{array}\right)\quad\textnormal{ for all }\zeta\in D_{1}.$ (12)
where $D_{1}=\{\zeta\in\mathcal{X}:\tau_{c}>0,\penalty 10000\ \tau_{g}=0\}$.
In Case (ii), a new input is applied to the system, a new output is sampled, and the timer $\tau_{c}$ resets to some point in the interval $\tau_{c,min},\tau_{c,max}]$, where $0<\tau_{c,min}\leq\tau_{c,max}$. This range of times represents indeterminacy in the amount of time that elapses between the application of successive inputs to the system. When the input changes, it is set equal to $z$. When $\tau_{c}$ reaches 0, the LTI system output is sampled and stored in $y_{s}$ based on Remark 3, which is held constant until the next sample. The states $x$ and $z$ do not change, and the jump map is
$G_{2}(\zeta)=\left(\begin{array}[]{c}x\\ z\\ Hu+d\\ z\\ {[\tau_{c,min},\tau_{c,max}]}\\ \tau_{g}\end{array}\right)\quad\textnormal{ for all }\zeta\in D_{2},$ (13)
where, $D_{2}\!=\!\{\zeta\in\mathcal{X}:\tau_{c}=0,\penalty 10000\ \tau_{g}>0\}$ and as described in Remark 3, we approximate the sampled output $y_{s}$ as $Hu+d$.
In Case (iii), we combine Cases (i) and (ii), and the system executes $G_{1}$ and then $G_{2}$ or $G_{2}$ and then $G_{1}$. The full jump map $G$ is defined as
@@
where $G_{3}(\zeta)=G_{1}(\zeta)\cup G_{2}(\zeta)$. Then, the full hybrid model of feedback optimization is
@@ IV-A Well-Posedness and Existence of Solutions
The high-level steps of the proof include: for a defined tangent cone $T_{C}$, applying Lemma 5, then prove that the only condition possible is that the solutions are complete, and concluded by ruling out Zeno behavior.
@@
Remark 4.
Even though Case (iii) in (14) introduces nondeterminism (whether $G_{1}$ or $G_{2}$ is first), this is an intended property that only applies two consecutive jumps (hence, Zeno is not possible).
@@ IV-B Algorithm Framework
We impose the following assumption about the computation of inputs and the value of the inputs applied.
Assumption 2.
Assumption 2 ensures that, when the system is properly initialized, at least $\ell$ gradient descent iterations are performed between any consecutive changes in the input. This condition is a mild form of timescale separation and it will be used later in Theorem 1.
Consider an initial condition $\phi({0},{0})=\nu\in\mathcal{X}$ that satisfies
$\tau_{c}(0,0)\in[\tau_{c,\min},\tau_{c,\max}],\tau_{g}(0,0)=\tau_{g,comp},\textnormal{ and }\\ z({0},{0})=u({0},{0}),$ (18)
and consider a solution $\phi$ to $\mathcal{H}_{FO}$ with those initial conditions. With regards to the state component $z$, we add a subscript to help denote how many gradient descent iterations there have been in computing the next value of the input $u$. More generally, we use $\alpha(i)$ to denote the number of iterates that are performed when computing the $(i+1)^{th}$ value of the input $u$. Assumption 2 implies that $\alpha(i)\geq\ell\geq 1$ for all $i\in\mathbb{N}$.
At the initial hybrid time (0,0)({0},{0}), the initial input to the system $u(0,0)$ is applied and held constant. The system then performs $\alpha(0)$ gradient descent iterations (which are $\alpha(0)$ Case (i) jumps) before the input to the LTI system is changed. When a Case (ii) jump occurs and is the $(\alpha(0)+1)^{th}$ jump, a new input is applied $u(t_{\alpha(0)+1},{\alpha(0)+1})$ where $\alpha(0)$ Case (i) jumps and one Case (ii) jump have occurred. When computing $u(t_{\alpha(0)+1},{\alpha(0)+1})$, the $k^{th}$ such iterate is denoted $z_{k}(t_{k},k)$ for any $k\in\{0,1,\cdots,\alpha(0)\}$. For computing $u(t_{\alpha(0)+1},{\alpha(0)+1})$, we denote the last iterate $\alpha(0)$ by $z_{\alpha(0)}(t_{\alpha(0)},{\alpha(0)})$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1)}=\\ z_{\alpha(p)}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha(p)+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha(p)+p),$ (21)
which indicates that the new value of the input is equal to the most recently computed optimization iterate. The initial iterate for computing the $(p+2)^{th}$ input is $z_{0}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1)$, and it is set equal to the same value as $u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1)$.